OpenStudy (anonymous):

Josh and his friends bought vanilla wafers for $4 per packet and chocolate wafers for $1 per packet at a carnival. They spent a total of $45 to buy a total of 27 packets of wafers of the two varieties. Part A: Write a system of equations that can be solved to find the number of packets of vanilla wafers and the number of packets of chocolate wafers that Josh and his friends bought at the carnival. Define the variables used in the equations. (5 points) Part B: How many packets of chocolate wafers and vanilla wafers did they buy? Explain how you got the answer and why you selected a partic

OpenStudy (anonymous):

@Michele_Laino can you help me on this problem also?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

yea thank you!

OpenStudy (michele_laino):

the system which models your problem is: \[\Large \left\{ \begin{gathered} x + y = 27 \hfill \\ 4x + y = 45 \hfill \\ \end{gathered} \right.\] where x is the number of packets of vanilla wafers, whereas y is the number of packets of chocolate wafers

OpenStudy (anonymous):

a) $4v + $1c = $45 v + c = 27 v = vanilla wafers c = chocolate wafers

OpenStudy (anonymous):

@Michele_Laino what does the x and y stand for. like v is vanilla and c is choclate?

OpenStudy (michele_laino):

yes! x=v and y=c

OpenStudy (anonymous):

ok wanted to be sure:)

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

so know what step is next?

OpenStudy (anonymous):

now*

OpenStudy (michele_laino):

we have to solve the first equation for y, for example, so we get: \[y = 27 - x\]

OpenStudy (michele_laino):

then we can substitute into the second equation: \[4x + \left( {27 - x} \right) = 45\]

OpenStudy (michele_laino):

please solve that equation for x

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

or 6 vanilla

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

b) solve by elimination: 1) 4v + c = 45 -1(v + c) = -1(27) 2) Distribute 4v + c = 45 -v - c = -27 3) Combine like terms: 3v =18 4) Solve for v: v = 6 5) plug v = 6 back into your original equation: 6 + c = 27 6) solve for c: c = 21 v=6 c=21

OpenStudy (michele_laino):

now we can substitute x=6, into this expression: \[\Large y = 27 - x = 27 - 6 = ...?\]

OpenStudy (michele_laino):

what do you get?

OpenStudy (anonymous):

21

OpenStudy (michele_laino):

ok! So, the solution of your system is: x=6, and y=21

OpenStudy (anonymous):

sweet. can you medal @thekefentse for me so he can get an award? ill medal you

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

than you:) also im probabley going to have mabe like two more questions if you can help me with them. ill post new and tag you

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Awesome! I hope i helped you to learn as well as @Michele_Laino

OpenStudy (anonymous):

@thekefentse ima post new question do you think you and @Michele_Laino can help on them if yall dont mind?

OpenStudy (anonymous):

I don't mind

OpenStudy (anonymous):

thank you. tag ya in a bit

OpenStudy (anonymous):

alright